OpenStudy (anonymous):

Austin and Kaitlyn Kojan invested$195,000 in a business venture if Kaitlyn invested 2 1/4 times as much as Austin invested how much money did Kaitlyn invest

OpenStudy (anonymous):

Ok so lets say Kaitlyn's investment is x and Austin's is y: We know:\[x+y=195000\] and that \[\frac{ 9 }{ 4 }y=x\] (2 1/4 equals 9/4)(Remember x is kaitlyn y is austin) Now we know that:\[x=\frac{ 9 }{ 4 }y\] So lets use that for x (kaitlyn's investment) in the original equation so: \[\frac{ 9 }{ 4 }y+y=195000\] Simplify:\[\frac{ 13 }{ 4 }y=195000\] Divide:\[y=60000\](Austin's investment) Now that we know y (Austin's investment) we can solve for x (kaitlyn's investment): \[x+60000=195000\] Subtract 60000: \[x=135000\](kaitlyn's investment) THE ANSWER IS: KAITLYN INVESTED 135,000 DOLLARS AND AUSTIN INVESTED 60,000 DOLLARS.